Institutt for fysikk NTNU

FY1002/TFY4160 Bølgefysikk Midtsemesterprøve fredag 10. oktober 2008 ca kl 12 – 14.

Fasit Versjon 0

Oppgave  A  B  C  D  Oppgave  A  B  C  D
1  x  14  x
2  x  15  x
3  x  16  x
4  x  17  x
5  x  18  x
6  x  19  x
7  x  20  x
8  x  21  x
9  x  22  x
10  x  23  x
11  x  24  x
12  x  25  x
13  x